(A) You are tasked with reporting the speeds for a road to the DOT. The speeds are shown in the table below, where for example 8 cars had a 30 mph speed, 12 cars a 33 mph speed, etc.

Speed Count
30 mph 8
33 mph 12
40 mph 17
42 mph 4
45 mph 9

What is the time-mean speed and space-mean speed for this traffic stream?

(B) You notice that 53% of the headways are 8 seconds or greater. What is the:

(i) Probability that exactly six vehicles come in the next minute?

(ii) Probability of three or less vehicles coming in the next minute?

(C) A roadway has a free-flow speed of 65 mph. When the road’s density is 120 veh/mi, the average space-mean speed is 45 mph. What is the road’s flow rate when the average space-mean speed is 40 mph?
Answer #1

Ans: A Time . mean speed I Svi qive t = 3088 + 33x12 +17840 +4x42 t q xus 8+12+17+4+9 Ve = 3718 vpb . Space mean speed (Vs)a) for trabfee stream. from green shield Model q=Kq v-/kilva RO = K; x 45-(K1x 452 = 45 kj-31.15 kj ..Ki= 120 = 8.66 veh/mi 1